# SpokenWOZ: A Large-Scale Speech-Text Benchmark for Spoken Task-Oriented Dialogue Agents

### v6

@@ A.9.1 Prompts for LLMs
| - Slot Name: hotel-pricerange; Slot Descrption: price budget of the hotel; Possible values: [’expensive’, ’cheap’, ’moderate’] |
| - Slot Name: hotel-type; Slot Descrption: type of the hotel; Possible values: [’guest house’, ’hotel’] |
| - Slot Name: hotel-parking; Slot Descrption: whether the hotel has parking; Possible values: [’no’, ’yes’] |
| - Slot Name: hotel-day; Slot Descrption: day of the hotel booking; Possible values: [’monday’, ’tuesday’, |
@@
| - Slot Name: hotel-people; Slot Descrption: number of people booking the hotel; Possible values: [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, |
@@
| - Slot Name: hotel-stay; Slot Descrption: length of stay at the hotel; Possible values: [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’7’, ’8’] |
| - Slot Name: hotel-internet; Slot Descrption: whether the hotel has the free internet; Possible values: [’no’, ’yes’] |
| - Slot Name: hotel-name; Slot Descrption: name of the hotel; Possible values: [] |
| - Slot Name: hotel-area; Slot Descrption: area of the hotel; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’] |
| - Slot Name: hotel-star; Slot Descrption: star of the hotel;Possible values: [’0’, ’1’, ’2’, ’3’, ’4’, ’5’] |
| - Slot Name: train-arriveby; Slot Descrption: the arrival time of the train, 24-hour standard time, e.g. 06:00, 18:30; |
@@
| - Slot Name: train-day; Slot Descrption: day of the train departure; Possible values: [’monday’, ’tuesday’, ’wednesday’, |
@@
| - Slot Name: train-people; Slot Descrption: number of people travelling by train; Possible values: [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, |
@@
| - Slot Name: train-leaveat; Slot Descrption: leaving time of the train, 24-hour standard time, e.g. 06:00, 18:30; |
@@
| - Slot Name: train-destination; Slot Descrption: destination of the train; Possible values: [’birmingham new street’, |
@@
| - Slot Name: train-departure; Slot Descrption: departure of the train; Possible values: [’birmingham new street’, |
@@
| - Slot Name: attraction-area; Slot Descrption: area of the attraction; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’] |
| - Slot Name: attraction-name; Slot Descrption: name of the attraction; Possible values: [] |
| - Slot Name: attraction-type; Slot Descrption: type of the attraction; Possible values: [’architecture’, ’boat’, ’cinema’, |
@@
| - Slot Name: restaurant-pricerange; Slot Descrption: price budget for the restaurant; Possible values: [’expensive’, ’cheap’, |
@@
| - Slot Name: restaurant-area; Slot Descrption: area of the restaurant; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’] |
| - Slot Name: restaurant-food; Slot Descrption: the cuisine of the restaurant; Possible values: [] |
| - Slot Name: restaurant-name; Slot Descrption: name of the restaurant; Possible values: [] |
| - Slot Name: restaurant-day; Slot Descrption: day of the restaurant booking; Possible values: [’monday’, ’tuesday’, |
@@
| - Slot Name: restaurant-people; Slot Descrption: number of people for the restaurant booking; Possible values: [’1’, ’2’, ’3’, |
@@
| - Slot Name: restaurant-time; Slot Descrption: time of the restaurant booking, 24-hour standard time, e.g. 06:00, 18:30; |
@@
| - Slot Name: hospital-department; Slot Descrption: department of the hospital; Possible values: [] |
| - Slot Name: taxi-leaveat; Slot Descrption: leaving time of taxi, 24-hour standard time, e.g. 06:00, 18:30; Possible values: [] |
| - Slot Name: taxi-destination; Slot Descrption: destination of taxi; Possible values: [] |
| - Slot Name: taxi-departure; Slot Descrption: departure location of taxi; Possible values: [] |
| - Slot Name: taxi-arriveby; Slot Descrption: arrival time of taxi, 24-hour standard time, e.g. 06:00, 18:30; Possible values: [] |
| - Slot Name: profile-name; Slot Descrption: the name of the user; Possible values: [] |
| - Slot Name: profile-email; Slot Descrption: the email of the user; Possible values: [] |
| - Slot Name: profile-idnumber; Slot Descrption: the idnumber of the user; Possible values: [] |
| - Slot Name: profile-phonenumber; Slot Descrption: the phonenumber of the user; Possible values: [] |
| - Slot Name: profile-platenumber; Slot Descrption: the platenumber of the user; Possible values: [] |

### v7

@@ A.9.1 Prompts for LLMs
- Slot Name: hotel-pricerange; Slot Descrption: price budget of the hotel; Possible values: [’expensive’, ’cheap’, ’moderate’] |
- Slot Name: hotel-type; Slot Descrption: type of the hotel; Possible values: [’guest house’, ’hotel’] |
- Slot Name: hotel-parking; Slot Descrption: whether the hotel has parking; Possible values: [’no’, ’yes’] |
- Slot Name: hotel-day; Slot Descrption: day of the hotel booking; Possible values: [’monday’, ’tuesday’, |
@@
- Slot Name: hotel-people; Slot Descrption: number of people booking the hotel; Possible values: [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, |
@@
- Slot Name: hotel-stay; Slot Descrption: length of stay at the hotel; Possible values: [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’7’, ’8’] |
- Slot Name: hotel-internet; Slot Descrption: whether the hotel has the free internet; Possible values: [’no’, ’yes’] |
- Slot Name: hotel-name; Slot Descrption: name of the hotel; Possible values: [] |
- Slot Name: hotel-area; Slot Descrption: area of the hotel; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’] |
- Slot Name: hotel-star; Slot Descrption: star of the hotel;Possible values: [’0’, ’1’, ’2’, ’3’, ’4’, ’5’] |
- Slot Name: train-arriveby; Slot Descrption: the arrival time of the train, 24-hour standard time, e.g. 06:00, 18:30; |
@@
- Slot Name: train-day; Slot Descrption: day of the train departure; Possible values: [’monday’, ’tuesday’, ’wednesday’, |
@@
- Slot Name: train-people; Slot Descrption: number of people travelling by train; Possible values: [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, |
@@
- Slot Name: train-leaveat; Slot Descrption: leaving time of the train, 24-hour standard time, e.g. 06:00, 18:30; |
@@
- Slot Name: train-destination; Slot Descrption: destination of the train; Possible values: [’birmingham new street’, |
@@
- Slot Name: train-departure; Slot Descrption: departure of the train; Possible values: [’birmingham new street’, |
@@
- Slot Name: attraction-area; Slot Descrption: area of the attraction; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’] |
- Slot Name: attraction-name; Slot Descrption: name of the attraction; Possible values: [] |
- Slot Name: attraction-type; Slot Descrption: type of the attraction; Possible values: [’architecture’, ’boat’, ’cinema’, |
@@
- Slot Name: restaurant-pricerange; Slot Descrption: price budget for the restaurant; Possible values: [’expensive’, ’cheap’, |
@@
- Slot Name: restaurant-area; Slot Descrption: area of the restaurant; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’] |
- Slot Name: restaurant-food; Slot Descrption: the cuisine of the restaurant; Possible values: [] |
- Slot Name: restaurant-name; Slot Descrption: name of the restaurant; Possible values: [] |
- Slot Name: restaurant-day; Slot Descrption: day of the restaurant booking; Possible values: [’monday’, ’tuesday’, |
@@
- Slot Name: restaurant-people; Slot Descrption: number of people for the restaurant booking; Possible values: [’1’, ’2’, ’3’, |
@@
- Slot Name: restaurant-time; Slot Descrption: time of the restaurant booking, 24-hour standard time, e.g. 06:00, 18:30; |
@@
- Slot Name: hospital-department; Slot Descrption: department of the hospital; Possible values: [] |
- Slot Name: taxi-leaveat; Slot Descrption: leaving time of taxi, 24-hour standard time, e.g. 06:00, 18:30; Possible values: [] |
- Slot Name: taxi-destination; Slot Descrption: destination of taxi; Possible values: [] |
- Slot Name: taxi-departure; Slot Descrption: departure location of taxi; Possible values: [] |
- Slot Name: taxi-arriveby; Slot Descrption: arrival time of taxi, 24-hour standard time, e.g. 06:00, 18:30; Possible values: [] |
- Slot Name: profile-name; Slot Descrption: the name of the user; Possible values: [] |
- Slot Name: profile-email; Slot Descrption: the email of the user; Possible values: [] |
- Slot Name: profile-idnumber; Slot Descrption: the idnumber of the user; Possible values: [] |
- Slot Name: profile-phonenumber; Slot Descrption: the phonenumber of the user; Possible values: [] |
- Slot Name: profile-platenumber; Slot Descrption: the platenumber of the user; Possible values: [] |